### v2

# Background Semantics Matter: Cross-Task Feature Exchange Network for Clustered Infrared Small Target Detection With Sky-Annotated Dataset

Abstract
Infrared small target detection poses unique challenges due to the scarcity of intrinsic target features and the abundance of similar background distractors. We argue that background semantics play a pivotal role in distinguishing visually similar objects for this task. To address this, we introduce a new task—-clustered infrared small target detection, and present DenseSIRST, a novel benchmark dataset that provides per-pixel semantic annotations for background regions, enabling the transition from sparse to dense target detection. Leveraging this dataset, we propose the Background-Aware Feature Exchange Network (BAFE-Net), which transforms the detection paradigm from a single task focused on the foreground to a multi-task architecture that jointly performs target detection and background semantic segmentation. BAFE-Net introduces a dynamic cross-task feature hard-exchange mechanism to embed target and background semantics between the two tasks. Furthermore, we propose the Background-Aware Gaussian Copy-Paste (BAG-CP) method, which selectively pastes small targets into sky regions during training, avoiding the creation of false alarm targets in complex non-sky backgrounds. Extensive experiments validate the effectiveness of BAG-CP and BAFE-Net in improving target detection accuracy while reducing false alarms. The DenseSIRST dataset, code, and trained models are available at https://github.com/GrokCV/BAFE-Net.
Index Terms:
Infrared Small Target Detection; Clustered Target Detection; Dataset; Feature Exchange; Semantic Segmentation
I Introduction
Figure 1: Illustration of the background semantics’ crucial role in infrared small target detection. The small target in the orange box represents a genuine target of interest against the sky background, while the small target in the blue box is a false alarm in the context of the building background. When considering only local regions, it is challenging to differentiate between clustered small targets and false alarms due to their similar appearances. By incorporating global contextual information, the distinction between real targets and false alarms becomes more apparent.
Infrared imaging has emerged as a crucial tool in various domains, particularly in scenarios where long-distance imaging is required in the absence of visible light, such as night-time operations or environments with limited visibility [1]. By capturing the thermal distribution of objects, infrared imaging provides critical information for applications such as unmanned system monitoring, maritime target search [2], and rescue missions [3]. However, in these applications, the targets of interest are often located at great distances, resulting in their appearance as extremely small-sized objects in the captured images. Consequently, infrared small target detection plays a pivotal role in accurately interpreting and extracting valuable insights from these images.
I-A Prior Works on Clustered Infrared Small Target Detection
While the potential of infrared imaging has been recognized across various domains, existing infrared small target detection techniques are predominantly tailored for the detection of sparse targets [4, 5], which proves to be inadequate when faced with complex application scenarios such as drone swarm defense, multi-target tracking systems, and extensive maritime search and rescue operations. Consequently, there is an urgent need for advanced technology capable of effectively identifying and tracking clustered infrared small targets.
As the distribution of small targets transitions from sparse to dense scenarios, existing infrared small target detection techniques face a multitude of challenges that drastically hinder their effectiveness. Firstly, methods that rely on the assumption of sparsely distributed targets become inapplicable in dense scenes, where targets are closely situated and may occlude one another, leading to the breakdown of single-target analysis strategies. Moreover, the reliance on local contrast features [6], a staple of traditional approaches, becomes unreliable as the targets are no longer set against a uniform background but are instead interspersed with other proximate objects. Furthermore, even state-of-the-art deep learning methods [7, 8, 9, 10] struggle when confronted with dense target scenarios where the delineation between targets and background blurs, significantly adding to the complexity of detection tasks. For instance, Dai et al. [11] introduced a detection transformer designed for arbitrarily oriented objects, yet its applicability to dense infrared small target settings remains limited due to the lack of intrinsic target features. Therefore, current methods are plagued by inefficiencies when operating in dense settings, frequently resulting in missed detections and high false alarm rates. This calls for new solutions tailored to overcome the intricacies of dense infrared small target detection.
While the rapid advancements in generic object detection have led to state-of-the-art methods like TOOD [12] and CrossKD [13] that address dense object detection, these techniques cannot be directly applied to the unique challenges posed by infrared small target detection. The specialized nature of this field presents several critical factors that hinder the direct transplantation of generic vision approaches:
  1. 1.
    Scarcity of Intrinsic Target Features: The foremost challenge stems from the paucity of intrinsic target features. Unlike generic objects, which boast rich visual cues such as appearance, texture, and structure, infrared small targets are minuscule in size, merely popping out as nondescript blobs in the image. This “visual poverty” renders recognition and localization methods based on target-specific characteristics largely ineffectual.
  2. 2.
    Abundance of Similar Background Distractors: Infrared imagery is often plagued by a multitude of background distractors that bear a striking resemblance to the targets of interest. These distractors, such as noise spots and hot spots, also manifest as blobs, making it nearly impossible to discriminate between them and genuine targets based on local visual comparisons. Liao et al. [14] tackled similar challenges in RGB-thermal salient object detection, where distortions like noise and motion blur compounded the complexity of separating salient objects from background clutter.
These challenges inevitably lead us to ponder a fundamental question: What truly distinguishes visually similar objects in single-frame infrared small target detection, assigning them different class labels? Unraveling this enigma holds the key to developing effective and robust algorithms for this task.
I-B Motivation
In this paper, we posit that the answer to the aforementioned question lies in the background semantics. As illustrated in Fig. 1, visually similar small objects at a distance can be classified differently based on their surrounding context. A small object in the sky background is likely to be a target of interest, while the same object against a building backdrop might be a false alarm. This stark contrast underscores the critical role that background semantics play in assigning meaningful labels to visually similar objects.
However, it is crucial to highlight a significant gap in the current landscape of infrared small target detection research. Despite the availability of several public datasets [15, 3, 16, 10], none of these datasets provide detailed semantic annotations for the background regions. To enable the explicit semantic modeling of backgrounds, a prerequisite is the creation of a dataset that includes background class annotations. Such a dataset would provide a solid foundation for researchers to develop algorithms that can leverage the rich contextual information present in the background regions.
To address the aforementioned challenges, this study presents DenseSIRST, a new open-source infrared target detection benchmark dataset. Compared to previous datasets in this field, DenseSIRST has implemented two significant upgrades. Firstly, the dataset transitions from a simple sparse target detection task to the precise detection of dense, clustered targets, substantially increasing the task’s difficulty and practical value. Secondly, DenseSIRST performs detailed per-pixel semantic segmentation of the background, dividing it into sky and non-sky categories, thereby providing preliminary support for explicit semantic modeling of the background.
Building upon this foundation, we propose the Background-Aware Feature Exchange Network (BAFE-Net) framework, which transforms the infrared small target detection paradigm from a single target detection task that solely focuses on the foreground to a multi-task architecture that jointly performs target detection and background semantic segmentation. BAFE-Net introduces a dynamic cross-task feature hard-exchange module in the target detection head and background segmentation head, which selectively embeds target semantics and background semantics between the segmentation and detection tasks through a feature adapter. To the best of our knowledge, this paper is the first to propose a cross-task feature hard-exchange mechanism. Benefiting from the explicit modeling of background semantics, BAFE-Net can better understand the contextual information in the scene, thereby significantly improving the accuracy of target detection and greatly reducing the false alarm rate.
Finally, by leveraging the fine-grained semantic annotations of the background, we propose a Background-Aware Gaussian Copy-Paste (BAG-CP) method. During the training phase, this method selectively pastes small targets into the sky region based on different background environments during data augmentation. This approach avoids pasting small targets into complex non-sky background regions, such as urban ground, which can produce highly similar false alarm targets.
In summary, our contributions can be categorized into THREE main aspects:
  1. 1.
    We propose a new task—-clustered infrared small target detection, and have released a new dataset, DenseSIRST, annotated with background semantic segmentation.
  2. 2.
    Utilizing DenseSIRST, we propose and demonstrate the effectiveness of the Background-Aware Gaussian Copy-Paste (BAG-CP) method, which outperforms the original Copy-Paste method that ignores background semantics.
  3. 3.
    We propose the BAFE-Net, which achieves a dynamic mechanism of mutual promotion between segmentation and detection tasks by implementing selective feature hard-exchange with a cross-task adapter, based on explicit background modeling.
II Related Work
II-A Infrared Small Target Detection
Infrared small target detection methods can be broadly classified into two main categories: model-driven and data-driven approaches. The model-driven methods, in particular, encompass three distinct subclasses, each employing unique strategies to capture and represent the relationship between targets and their surrounding backgrounds.
  1. 1.
    Background Estimation Methods: By leveraging the local consistency of backgrounds, these methods construct filters to enhance target saliency. However, they falter when dealing with extremely weak targets, leading to missed detections and a high sensitivity to hyperparameters. The limitation stems from insufficient semantic understanding of the features [17], which results in poor performance in complex backgrounds and increased false alarms.
  2. 2.
    Human Visual System (HVS) Methods: Techniques such as local contrast measure (LCM) [6] and multi-scale patch-based contrast measure (MPCM) [18] attempt to emulate the high spatial local contrast sensitivity of the human eye. Despite their intuitive appeal, these methods struggle with very small or blurred targets, mainly due to their inability to generate sufficient contrast.
  3. 3.
    Low-Rank and Sparse Decomposition (LRSD) Methods: This category includes approaches like the infrared patch-image (IPI) model [4] and reweighted infrared patch-tensor (RIPT) model [5]. These methods have gained popularity due to their superior background suppression capabilities compared to previous techniques. However, they severely suffer from the challenge of distinguishing background distractors from targets due to the similar sparse characteristic.
Despite their contributions, these methods suffer from several limitations, including poor performance in complex backgrounds, high false alarm rates, and sensitivity to hyperparameters, which we attribute to the lack of feature semantics.
Transitioning from model-driven methods, data-driven approaches have revolutionized the field by leveraging deep networks to automatically learn features relevant for detecting small targets. Recent advancements have also focused on implementing these deep learning-based methods efficiently, such as the work on high-accuracy and real-time two-stage small object detection on FPGA [19]. Methods such as asymmetric context modulation (ACM) [15], dense nested attention networks (DNANet) [16], and gated shaped trans U-Net (GSTUnet) [9] have shown significant promise. These methods concentrate on integrating the features of infrared small targets with neural networks, utilizing deeper neural networks or more robust attention modules. Similar advancements in applying deep learning to infrared imaging have been seen in related fields, such as the work on infrared-guided eye localization for autostereoscopic displays [20]. The advantages of data-driven methods lie in their ability to automatically extract features and focus on multi-scale feature interaction, which can be seen as a form of implicit modeling of the relationship between targets and background. Similar approaches have been applied in other domains, such as cloud detection, where multi-level information fusion with edge information injection has been used to improve detection performance [21]. In a related field, transformer-based models have been employed for spatio-temporal unsupervised traffic anomaly detection in aerial videos [22], demonstrating the versatility of these advanced deep learning techniques in handling complex visual data with small objects or anomalies.
While effective to a degree, these methods treat all backgrounds as a single class, ignoring the significant impact of different background semantics on distinguishing between true targets and false alarms with similar appearances. Moreover, despite not fully depending on sparse priors or local contrast priors like traditional methods, some approaches still leverage this prior knowledge to enhance detection performance [23]. Their performance is greatly affected in dense scenes with multiple adjacent targets, where the assumption of local contrast between targets and their neighboring background no longer holds. This fundamental shift in target distribution necessitates the development of approaches that can effectively handle the complex interactions and dependencies among multiple targets in close proximity.
In contrast to previous studies, our work explores a new detection paradigm for infrared small targets:
  1. 1.
    Unexplored Challenge for Clustered Targets: Unlike traditional approaches that focus on sparse targets, we take on the more complex task of detecting clustered targets, introducing a new challenge in this field. We further facilitate the research by providing DenseSIRST, a dedicated dataset created specifically for this novel task.
  2. 2.
    Explicit Semantic Modeling: In contrast to previous studies that treat backgrounds as a homogeneous class, we leverage the DenseSIRST dataset’s semantic annotations for background modeling. Our proposed BAFE-Net thereby gains an enhanced understanding of the target’s background, leading to improved detection accuracy.
  3. 3.
    Introduction of Cross-Task Feature Exchange: Our dynamic cross-task feature exchange module brings together background semantics and target features through a feature adapter, a strategy not found in existing work. This approach enables a more precise differentiation between targets and their backgrounds.
II-B Data Augmentation for Object Detection
Data augmentation is pivotal in enhancing the performance and generalization of object detection models. By diversifying training data through various transformations such as flipping, rotation, scaling, and cropping, models learn to recognize and generalize across a broader spectrum of object variations. However, while these basic techniques are widely used, they often fall short in complex detection scenarios, necessitating more sophisticated methods.
To address the limitations of traditional methods, advanced augmentation techniques have been developed, which introduce more nuanced transformations that better preserve and highlight essential features of the objects while still enhancing the dataset’s diversity. For instance, CutMix [24] improves upon the CutOut [25] method by overlaying a segment of one image onto another, blending labels proportionally to the area of the patch, thereby maintaining more contextual integrity. Attentive-CutMix [26] refines this approach by selecting regions based on feature importance, thus preserving significant target details. SnapMix [27] and SaliencyMix [28] adjust the blending process based on saliency or importance maps, aiming to preserve crucial features while enhancing background complexity. TransMix [29] integrates attention mechanisms to better manage the blending of features, focusing on maintaining important target characteristics during augmentation. StyleMix [30] and PuzzleMix [31] introduce style and structural considerations into the mix, allowing for more complex transformations that consider both content and style elements of the images, potentially offering more robust augmentation strategies.
Despite these advancements, these methods primarily assume the presence of larger, easily distinguishable targets and do not specifically cater to the unique challenges posed by clustered infrared small target detection. None of these methods are explicitly designed to handle the high-density and occlusion-prone nature of infrared small targets. This gap highlights the need for augmentation strategies that directly address the intricacies of detecting small, closely-spaced targets against complex backgrounds.
Although our BAG-CP aligns with the broader goal of enhancing infrared small target detection, it stands out in its unique focus and strategies:
  1. 1.
    Pioneering Data Augmentation for Clustered Infrared Small Target Detection: Uniquely, our method pioneers a data augmentation technique specifically tailored for clustered infrared small target detection. It takes into account the local contrast prior and enables dynamic random variations in local contrast, a perspective not considered in existing techniques.
  2. 2.
    Background-Aware Target Placement: Our BAG-CP approach diverges from the commonly used random or fixed-position copy-paste tactics. Instead, it integrates background semantics to guide target placement, thereby skirting potential category conflicts induced by inconsistent background contextual semantics.
II-C Multi-Task Learning and Cross-Task Mechanism
Multi-task learning (MTL) is a pivotal strategy in computer vision that enhances model performance and generalization by jointly learning shared representations from related tasks, such as detection and segmentation. Takikawa et al. [32] introduced a two-stream CNN architecture for semantic segmentation that explicitly wires shape information as a separate processing branch, using gates to connect the intermediate layers of the two streams. Misra et al. [33] proposed a new sharing unit called “cross-stitch” unit, which combines activations from multiple networks and can be trained end-to-end to learn an optimal combination of shared and task-specific representations. Zhang et al. [10] developed ISNet, which employs a Taylor finite difference inspired edge block and two-orientation attention aggregation block to enhance edge information and capture shape characteristics of infrared small targets.
However, these methods may not be well-suited for clustered infrared small target detection, as they often rely on edge detection or surface normal estimation, which assume larger targets. In real-world scenarios, infrared small targets typically consist of only 1 to 9 pixels and are subject to the diffusion effect, lacking clear edges and falling far below the conventional definitions of small targets, i.e., $32\times 32$ pixels in general object detection.
In this context, cross-task mechanisms that enable semantic exchange between different tasks are more relevant to our work. These methods enhance MTL by facilitating feature exchange or distillation across different tasks, allowing each task to benefit from complementary signals of others. Methods like prediction-and-distillation network utilize attention-based mechanisms to distill multi-modal features [34], while MTI-Net extends this concept with a multi-scale approach better suited for handling diverse task demands [35].
While our research shares the underpinnings of multi-task learning and cross-task mechanisms, we pioneer a new approach to feature fusion. Instead of following the well-trodden path of soft fusion found in attention-based methods, our work introduces an innovative protocol–-a dynamic hard-exchange mechanism. To the best of our knowledge, this is the first work to explore the potential of dynamic hard-exchange of features between tasks.
III The DenseSIRST Dataset
Figure 2: Illustration of selected images from our proposed DenseSIRST dataset. Each image pair consists of a simulated dense small target image (top) and its corresponding sky-segmented version (bottom). These images highlight the challenges posed by densely clustered targets within diverse backgrounds.
III-A Motivation and Rationale
The detection of small, low-contrast targets against complex backgrounds in infrared imagery poses a significant challenge. Existing datasets, though instrumental in progressing the field, exhibit critical deficiencies. As shown in Table I, well-regarded datasets such as SIRST V1 [15], SIRST V2 [3], IRSTD1K [10], and SIRSTAUG [36] predominantly feature sparsely distributed targets and lack per-pixel background annotations. This stands in stark contrast to the dense, diminutive targets frequently encountered in real-world scenarios.
To bridge this gap, we present DenseSIRST: a meticulously annotated dataset of infrared images teeming with densely clustered, small targets, which is a more realistic and challenging benchmark for algorithm training and evaluation. Fig. 2 showcases a selection of images from the DenseSIRST dataset, illustrating the dense distribution of small targets and the corresponding sky segmentation annotations. These annotations provide valuable contextual information for developing and evaluating detection algorithms.
TABLE I: Comparison of the DenseSIRST dataset with other popular datasets used in infrared small target detection.
Datasets Image Type Annotation Type Image Number Target Number Average Target Area Sparse or Clustered Background Semantic Annotation
SIRST V1 [15] Real Pixel 427 533 23 Sparse
SIRST V2 [3] Real Pixel + BBox + Point 1024 648 24 Sparse
IRSTD1K [10] Real Pixel 1001 1495 38 Sparse
SIRSTAUG [36] Synthetic Pixel 8525 9278 88 Sparse
DenseSIRST (Ours) Real + Synthetic Pixel + BBox + Point 1024 13655 6 Cluster
Figure 3: Statistical characteristics of our DenseSIRST dataset, underscoring the detection challenges it presents. (a) The distribution of target sizes, with the circle sizes in the scatter plot indicating the prevalence of each size category. (b) The distribution of local contrasts for small targets in the dataset, demonstrating that the dataset encompasses a broad range of contrasts, down to a minimum of 0.15. (c) The brightness distribution of small targets in the dataset, highlighting the wide spectrum from dark to light targets this dataset offers for detection tasks.
III-B Target Synthesis and Composition
The DenseSIRST dataset, a blend of real and synthetic infrared images, provides a challenging and dense environment for training robust detection algorithms. We curated this dataset from diverse real-world scenarios, supplemented with synthetic targets designed for physical plausibility and visual realism.
The dataset leverages our SIRST V2 dataset, creating a broad library of infrared small targets. Each image is meticulously designed to emulate the densely populated surveillance scenarios. Specifically, one to three “dense areas” are defined randomly within the less cluttered sky regions of each base image. These areas, each measuring $20\times 20$ pixels, are populated with 8 to 12 targets selected randomly from the target library. The selected targets are resized to dimensions within $5\times 5$ pixels to maintain scale variability and pasted with a spacing of 1 to 2 pixels between them to simulate realistic clustering.
To ensure a smooth integration of targets and their backgrounds, we employ a Gaussian-weighted copy-paste technique, an evolution of the concept from mixup. This modification is governed by the following equation:
$T_{new}=T_{back}+B_{box}\times\lambda\times G\ ,\lambda\in[0.5,1],$ (1)
where $T_{new}$ denotes the adjusted target appearance, while $T_{back}$ is the original background at the target’s location. Different from mixup, its weight is set to $1$ for enhancing the reality of the synthesized target. $B_{box}$ represents the original target appearance. The scalar $\lambda$, which adjusts the target’s brightness, is randomized to diversify the brightness and contrast of targets, thus enriching the sample space.
The matrix $G$ denotes a Gaussian blur, employed during target pasting, and is defined as:
$G(x,y)=e^{-\frac{{(x-\rho_{x}w)^{2}}}{{2\sigma_{x}^{2}}}-\frac{{(y-\rho_{y}h)^{2}}}{{2\sigma_{y}^{2}}}},$ (2)
where the relative offsets $\rho_{x}$ and $\rho_{y}$ determine the Gaussian distribution center along the $x$ and $y$ axes, randomly selected within the range $[0,0.2]$. These values, when multiplied by the target bounding box’s width $w$ and height $h$, yield the actual pixel offsets, ensuring appropriate scaling with the target size. The Gaussian distribution’s spread is controlled by the standard deviations $\sigma_{x}$ and $\sigma_{y}$ along the $x$ and $y$ axes, randomized within $[0.3,0.6]$. This randomness guarantees varying degrees of edge blurring among the simulated targets, reinforcing their realistic appearance. To further enhance variability, each target undergoes a rotational transformation. The Gaussian matrix’s rotation $\theta$ is applied as:
$\begin{bmatrix}x^{\prime}\\ y^{\prime}\end{bmatrix}=\begin{bmatrix}\cos(\theta)&-\sin(\theta)\\ \sin(\theta)&\cos(\theta)\end{bmatrix}\begin{bmatrix}x-\mu_{x}\\ y-\mu_{y}\end{bmatrix},$ (3)
where $\theta$ is randomly selected from the interval $\left[-90^{\circ},90^{\circ}\right]$.
III-C Dataset Description and Characteristics
III-C1 Dataset Composition
The DenseSIRST dataset comprises 1024 infrared images, with a total of 13,655 densely clustered small targets. The images in the DenseSIRST dataset cover a wide range of realistic scenarios, including urban areas, mountainous regions, maritime environments, and cloudy scenes. This diversity enables the development and assessment of detection algorithms that can generalize well to various real-world applications. The division of training, validation, and test set is aligned with the SIRST V2 dataset.
III-C2 Annotations and Labels
In the DenseSIRST dataset, each image is meticulously annotated with three types of labels: pixel-level masks, bounding boxes, and point annotations. The pixel-level masks provide precise segmentation of each target, enabling the evaluation of algorithms at a fine-grained level. Bounding box annotations encapsulate each target with a rectangular box, facilitating the assessment of detection performance using metrics such as precision, recall, and F1 score. Point annotations mark the center of each target, allowing for the evaluation of localization accuracy.
III-C3 Statistical Characteristics
The DenseSIRST dataset is meticulously curated to provide a comprehensive benchmark for clustered infrared small target detection. This part offers an analysis of the dataset’s key characteristics.
Target Size Distribution: Most targets, as indicated in Fig. 3 (b), are extremely small, predominantly below $5\times 5$ pixels. The $3\times 3$ pixels size is the most frequent, aligning with the characteristics of small target detection in infrared imagery.
Local Contrast Distribution: The 90% targets have local contrast below $2$, as shown in Fig. 3 (c). This low contrast is indicative of the inherent difficulty in detecting such targets against cluttered backgrounds. Moreover, the presence of nearby targets within dense clusters can inadvertently be included in the background calculation, leading to an underestimation of local contrast and further complicating detection.
Target Brightness Distribution: Fig. 3 (d) reveals that only a minority of targets are the brightest points in their respective images. This characteristic emphasizes the inadequacy of brightness as a sole distinguishing feature for target detection.
In conclusion, the DenseSIRST dataset, with its wide range of image dimensions, prevalence of small target sizes, and low local contrast, provides a realistic and challenging benchmark for infrared small target detection algorithms.
IV Background-Aware Feature Exchange Network
IV-A Overall Architecture and Training Pipeline
Figure 4: Schematic diagram of the proposed BAFE-Net network. BAFE-Net contains three modules: ResNet, FPN and BAFE-Head.
Figure 5: First, a dense area is randomly selected from the sky area of the image, and then several small targets are randomly selected from the small target library. After determining the appropriate pasting position, the Gaussian technique is used to integrate the small targets into the new background.
The overall architecture of BAFE-Net, which adopts the fully convolutional one-stage object detector (FCOS) [37] as its host network, is illustrated in Fig. 4. In addition to the original detection head present in the baseline, BAFE-Net introduces a segmentation head within the BAFE-Head module. The primary purpose of incorporating this segmentation head is to enable explicit modeling of background semantics, which is crucial for enhancing the network’s ability to distinguish between real targets and background clutter in infrared scenes. By jointly optimizing the detection and segmentation tasks, BAFE-Net learns to capture both object-level and pixel-level contextual information. Furthermore, BAFE-Net introduces a dynamic cross-task feature hard-exchange mechanism to facilitate the bidirectional flow of information between the detection and segmentation branches. This dynamic exchange allows the network to adaptively leverage the most relevant features from each task, leading to improved detection performance.
To address the challenge of limited target samples, BAFE-Net incorporates a novel training methodology termed Background-Aware Gaussian Copy-Paste (BAG-CP). This technique strategically places synthetic infrared small targets into real scene backgrounds by considering semantic context, ensuring that the pasted targets are only placed in sky regions. The use of Gaussian smoothing at the edges of the pasted targets ensures seamless integration, avoiding artificial boundary artifacts that could mislead the learning algorithm. By augmenting the training dataset with realistically integrated synthetic targets, BAG-CP allows BAFE-Net to train on a dataset that closely mimics operational scenarios, where small targets often appear against complex and cluttered backgrounds. This context-aware synthesis and integration approach enables BAFE-Net to effectively recognize and handle these challenging conditions, ultimately setting a new standard for performance in infrared small target detection.
IV-B Dynamic Cross-Task Feature Exchange Head
For infrared small target detection, the scarcity of intrinsic target features and the abundance of similar background distractors pose significant challenges. To address this, we propose the Background-Aware Feature Exchange Network (BAFE-Net), which transforms the detection paradigm from a single task focused on the foreground to a multi-task architecture that jointly performs target detection and background semantic segmentation. The motivation behind this cross-task feature exchange is twofold:
  1. 1.
    Complementary Nature of Tasks: Although target detection and background segmentation have different objectives, they share a common goal of understanding the scene. By exchanging features between these tasks, each task can benefit from the complementary information learned by the other, leading to a more comprehensive understanding of the image.
  2. 2.
    Synergistic Relationship: Target detection focuses on identifying the presence and location of small targets, while background segmentation emphasizes the contextual information and local details of the scene. By sharing features between these tasks, target detection can leverage the semantic understanding of the background to better distinguish targets from distractors, while background segmentation can utilize the localized target information to refine its segmentation boundaries.
IV-B1 Dynamic Channel Selection
To further optimize cross-task feature exchange, we propose a dynamic selection mechanism that determines the most relevant feature channels for exchange based on their importance to task performance. The importance of each channel is estimated using an attention mechanism, which assigns weights to channels according to their contribution.
Given an input feature map $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$, where $C$, $H$, and $W$ denote the number of channels, height, and width of the feature map, respectively, we first compute a channel-wise attention weight vector $\mathbf{m}\in\mathbb{R}^{C}$ by a mechanism inspired by the design of our previous work Large Selective Kernel Network (LSKNet) [38]. This is achieved by applying a series of operations to $\mathbf{X}$, including local feature extraction, channel-wise statistics computation, and weighted fusion. The process can be summarized as follows:
$\mathbf{m}$ $=\sigma\left(\mathbf{X}\odot\left(\mathcal{F}(\mathbf{X})\odot\sigma\left(\mathcal{S}\left(\text{avg}(\mathbf{X})\oplus\max(\mathbf{X})\right)\right)\right)\right),$ (4)
where $\odot$ denotes element-wise multiplication, $\oplus$ denotes the concatenation operation, and $\sigma$ is the Sigmoid function. $\max$ and avg are the maximum and average pooled spatial feature descriptors. $\mathcal{S}$ is a convolution operation that reduces the dimensionality of concatenated features. $\mathcal{F}$ is also a convolution operation applied to $\mathbf{X}$ , extracting local features.$\mathbf{m}\in\mathbb{R}^{C}$ is a one-dimensional tensor, and $C$ is the number of channels. Each element $\mathbf{m}$ represents the attention logits for the $i$-th channel.
For each sample, the top $K$ channels with the highest weights are selected, where $K$ is determined as a fraction $p$ of the total number of channels $C$:
$K$ $=C\times p,$ (5)
${V}_{\text{topk}},{I}_{\text{topk}}$ $=\mathrm{TopK}({\mathbf{m}},K),$
where ${V}_{\text{topk}}\in\mathbb{R}^{K}$ represents the top $K$ largest weight values for each sample. ${I}_{\text{topk}}\in\mathbb{R}^{K}$ represents the corresponding channel indices for these weights. This process dynamically selects the channels for exchange, improving the effectiveness of cross-task feature sharing.
IV-B2 Cross-Task Feature Hard-Exchange
Within BAFE-Net, we introduce a cross-task feature hard-exchange mechanism to embed target and background semantics between the two tasks. Feature hard-exchange allows for the direct transfer of learned features from one task to another. By swapping entire feature channels, the receiving task has access to the explicit representation learned by the other task, facilitating a more direct integration of complementary information.
To perform feature exchange, let $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$ be the input tensors from the target detection and background segmentation tasks, respectively, each with a shape of $(C,H,W)$, where $C$ is the number of channels, and $H$ and $W$ are the height and width. We define an exchange mask $\mathbf{e}\in\mathbb{R}^{C}$:
$\mathbf{e}[i]=\begin{cases}1,&\mathrm{if}\ i\in I_{\text{topk}}\\ 0,&\mathrm{otherwise}\end{cases},$ (6)
where $i$ represents the channel index. The exchange mask $\mathbf{e}$ is then expanded to the shape of $(C,H,W)$:
$\mathbf{E}_{d}=\mathbf{e}\otimes\mathbf{1}_{H\times W},$ (7)
where $\otimes$ denotes the expansion operation, and $\mathbf{1}_{H\times W}$ is a tensor of ones with a shape of $(H,W)$. $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$ can obtain $\mathbf{E}_{d1}$ and $\mathbf{E}_{d2}$ respectively through Eq. (4) to Eq. (7). The output tensors $\mathbf{Y}_{1}$ and $\mathbf{Y}_{2}$ after the hard feature exchange are computed as:
$\mathbf{Y}_{1}=\mathbf{X}_{1}\odot(\mathbf{1}-\mathbf{E}_{d1})+\mathcal{A}\left(\mathbf{X}_{2}\odot\mathbf{E}_{d2}\right),$ (8)
$\mathbf{Y}_{2}=\mathbf{X}_{2}\odot(\mathbf{1}-\mathbf{E}_{d2})+\mathcal{A}\left(\mathbf{X}_{1}\odot\mathbf{E}_{d1}\right),$ (9)
where $\odot$ denotes element-wise multiplication, $\mathbf{Y}_{1}$ and $\mathbf{Y}_{2}$ represent the output features, and $\mathbf{E}_{d1}$ and $\mathbf{E}_{d2}$ are the dynamic masks. To address the cross-modal nature of features, we introduce an adapter module $\mathcal{A}$, implemented as a multi-layer perceptron (MLP), which projects features into a common space before exchange. This design enables effective cross-modal interaction while preserving modality-specific characteristics.
IV-C Background-Aware Gaussian Copy-Paste
Infrared small target detection is inherently challenged by the paucity of intrinsic visual features in targets and the interference from similar background elements. Traditional data augmentation strategies, such as random cropping or flipping, are not effective in addressing these unique challenges. Consequently, we propose a novel data augmentation technique, Background-Aware Gaussian Copy-Paste (BAG-CP), which capitalizes on contextual information to enhance the training process. BAG-CP selectively augments training images by pasting infrared small targets into semantically relevant background areas, generating more realistic and diverse training samples. The fundamental idea is to bolster the detector’s ability to discern real targets from background noise through the seamless blending of pasted targets into their new environment. As illustrated in Fig. 5, The BAG-CP process unfolds as follows:
Prerequisite: Semantic Annotation of Background: Before performing the copy-paste operation, it is crucial to have the image segmented into different semantic regions. The DenseSIRST dataset includes annotated background categories, providing detailed background masks that separate sky from non-sky regions.
Step 1: Selection and Extraction of Infrared Small Targets: A library of small targets is created during the generation of the DenseSIRST dataset. A random selection of small targets from this library is made for pasting into the current training image.
Step 2: Gaussian Smoothing of Edges: Prior to pasting, the selected targets are subjected to Gaussian smoothing. This step ensures a smooth transition around the pasted targets, minimizing any conspicuous artifacts that might disrupt the learning process.
Step 3: Context-Aware Target Pasting Based on Background Semantics: The core of BAG-CP lies in context-aware placement. Using the semantic annotation of the background, we classify it into sky and non-sky regions. When choosing where to paste the Gaussian-smoothed targets, we only select locations within the sky regions. This approach mimics real-world scenarios where infrared small targets, especially aerial targets, are more likely to appear against a sky background.
During the pasting process, 1-3 regions are selected within the image to form clusters of small targets. Each region is 20 × 20 pixels in size and is populated with 8 to 12 targets randomly chosen from the target library. The selected targets are resized to within 5 × 5 pixels to maintain proportional variability and are pasted with a spacing of 1 to 2 pixels to simulate realistic clustering. The Gaussian-weighted edge blurring technique used follows Eq. (1), Eq. (2), and Eq. (3).
V Experiments
In this section, we outline a series of experimental evaluations designed to validate the effectiveness of the Background-Aware Gaussian Copy-Paste (BAG-CP) method and the Background-Aware Feature Exchange Network (BAFE-Net) proposed in this study. Additionally, we conduct comparative analyses with state-of-the-art methods in infrared small target detection, using both quantitative metrics and qualitative assessments. We address the following pivotal questions in our experimental analysis:
  1. 1.
    Ablation Study on BAG-CP. We compare the performance of various copy-paste strategies to demonstrate the superiority of our BAG-CP method.
  2. 2.
    Effectiveness of Background-Aware Feature Exchange. We evaluate the performance of BAFE-Net by comparing different configurations to highlight the contributions of each component in BAFE-Net.
  3. 3.
    Feature Exchange Mechanism Analysis. We investigate the impact of different feature exchange strategies on the performance of BAFE-Net to determine the optimal approach for exchanging features between the target detection head and the background segmentation head.
  4. 4.
    Impact of Backbone Depth: We examine the influence of varying backbone depths on the performance of the BAFE-Net. The primary aim is to establish the optimal depth that maximizes target detection while balancing computational efficiency.
  5. 5.
    Comparative Analysis with State-of-the-Art Methods. We conduct a comprehensive comparison between our method and other state-of-the-art methods on the DenseSIRST dataset.
V-A Experimental Settings
1) Dataset: During the experimental phase of our research, we conducted tests on the proposed dataset DenseSIRST.
2) Implementation Details: Considering the image sizes in the dataset and computational overhead, for methods within the DeepInfrared toolkit, we set the range of the long edge of training and testing images to [512, 512] while maintaining aspect ratio to avoid image distortion. We choose Focal Loss as the loss function of the classification branch, IoU Loss as the loss function of the regression branch, and Cross Entropy Loss as the loss function of the segmentation branch. We choose DAdaptAdam as the optimizer, learning rate is set to 1.0, and weight decay is set to 0.05. Considering training and convergence time, we trained for 20 epochs within the DeepInfrared framework.
3) Evaluation Metrics:We used mean Average Precision (mAP) with IoU thresholds of 0.5 ($mAP_{07}$) and 0.75 ($mAP_{12}$), and Recall as evaluation metrics for our experiment.
TABLE II: Comparison with Other State-of-the-art methods on DenseSIRST.
Method Backbone DenseSIRST
FLOPs $\downarrow$ Params $\downarrow$ mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
              One-stage
SSD [39] 87.552G 23.746M 0.211 0.421 0.178 0.424
RetinaNet [40] ResNet50 52.203G 36.330M 0.114 0.510 0.086 0.523
YOLOv3 [41] 50.002G 61.949M 0.233 0.424 0.207 0.413
CenterNet [42] ResNet50 50.278G 32.111M 0.138 0.316 0.124 0.317
FCOS [37] ResNet50 50.291G 32.113M 0.232 0.315 0.204 0.324
ATSS [43] ResNet50 51.504G 32.113M 0.248 0.327 0.202 0.326
CentripetalNet [44] 0.491T 0.206G 0.244 0.259 0.201 0.244
AutoAssign [45] ResNet50 50.555G 36.244M 0.255 0.354 0.180 0.314
GFL [46] ResNet50 52.296G 32.258M 0.264 0.367 0.230 0.317
PAA [47] ResNet50 51.504G 32.113M 0.255 0.545 0.228 0.551
VFNet [48] ResNet50 48.317G 32.709M 0.253 0.336 0.214 0.336
PVT-T [49] 41.623G 21.325M 0.109 0.481 0.093 0.501
YOLOF [50] ResNet50 25.076G 42.339M 0.091 0.009 0.002 0.009
YOLOX [51] 8.578G 8.968M 0.210 0.341 0.180 0.331
TOOD [12] ResNet50 50.456G 32.018M 0.256 0.355 0.226 0.342
DyHead [52] ResNet50 27.866G 38.890M 0.249 0.335 0.189 0.328
DDOD [53] ResNet50 46.514G 32.378M 0.253 0.335 0.230 0.351
RTMDet [54] 51.278G 52.316M 0.229 0.349 0.212 0.350
EfficientDet [55] 34.686G 18.320M 0.146 0.464 0.094 0.517
Two-stage
Faster R-CNN [56] ResNet50 0.759T 33.035M 0.091 0.022 0.015 0.029
Cascade R-CNN [57] ResNet50 90.978G 69.152M 0.136 0.188 0.139 0.194
Grid R-CNN [58] ResNet50 0.177T 64.467M 0.156 0.122 0.104 0.190
Libra R-CNN [59] ResNet50 63.990G 41.611M 0.141 0.142 0.085 0.120
TridentNet [60] ResNet50 0.759T 33.035M 0.091 0.009 0.014 0.021
SABL [61] ResNet50 0.125T 42.213M 0.124 0.104 0.104 0.171
Dynamic R-CNN [62] ResNet50 63.179G 41.348M 0.184 0.235 0.111 0.190
End2End
DETR [63] ResNet50 24.940G 41.555M 0.000 0.000 0.000 0.000
Sparse R-CNN [64] ResNet50 45.274G 0.106G 0.183 0.572 0.154 0.614
Deformable DETR [65] ResNet50 51.772G 40.099M 0.024 0.016 0.018 0.197
Conditional DETR [66] ResNet50 27.143G 40.297M 0.000 0.000 0.000 0.001
DAB-DETR [67] ResNet50 28.939G 43.702M 0.005 0.054 0.000 0.001
$\star$ BAFE-Net (Ours) ResNet18 57.654G 22.31M 0.283 0.335 0.233 0.325
$\star$ BAFE-Net (Ours) ResNet50 71.639G 35.626M 0.274 0.342 0.248 0.338
Figure 6: Feature visualization comparison. Compared with other methods, the proposed BAFE-Net shows superior performance with higher detection of small objects and fewer false alarms.
V-B Comparison with State-of-the-Arts
We benchmarked our BAFE-Net against current leading or state-of-the-art methods on the proposed DenseSIRST dataset, using metrics such as $mAP_{07}$, $mAP_{12}$, and Recall. The experimental results are presented in Table II.
End-to-end (End2End) methods, while theoretically appealing due to their NMS-free style, demonstrated suboptimal performance. This can be attributed to their inherent requirement for extensive training data, akin to COCO-level volumes, for effective operation. The relatively limited availability of infrared small target detection datasets inherently handicaps such methods, resulting in their diminished performance.
Next, we observed that two-stage methods underperformed compared to one-stage methods. Two-stage methods generate candidate regions based on deeper features, which are then subjected to classification and bounding box regression. However, the DenseSIRST dataset, characterized by small and densely clustered infrared targets, poses a challenge for such methods, leading to their suboptimal performance.
In contrast, one-stage methods, which directly regress bounding boxes and categories on the feature pyramids, demonstrated superior performance. These methods are not only faster and more computationally efficient, but also achieved higher mAP and recall, indicating their overall effectiveness. Significantly, our BAFE-Net method outperformed all other one-stage methods. The superior performance of BAFE-Net validates the efficacy of our approach in accurately detecting small infrared targets amidst dense conditions.
TABLE III: Quantitative ablation study evaluates the impact of the Copy-Paste technique, clustering of pasted objects, Gaussian module, and Background-Aware pasting within the data augmentation module of BAFE-Net
Strategy Module DenseSIRST
Copy-Paste Cluster Gaussian Background-Aware mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
  (a) 0.225 0.311 0.180 0.314
(b) $\checkmark$ 0.236 0.322 0.203 0.303
(c) $\checkmark$ $\checkmark$ 0.242 0.329 0.205 0.335
(d) $\checkmark$ $\checkmark$ 0.245 0.321 0.211 0.338
(e) $\checkmark$ $\checkmark$ $\checkmark$ 0.250 0.319 0.215 0.315
(f) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.258 0.340 0.220 0.340
TABLE IV: Quantitative ablation study evaluates the impact of the Background-Aware segmentation, Feature Exchange module, and adapter module within the Background-Aware Feature Exchange Network (BAFE-Net)
Strategy Module DenseSIRST
Seg Head Feature Exchange Adapter mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
  (a) 0.225 0.311 0.180 0.314
(b) $\checkmark$ 0.246 0.324 0.183 0.304
(c) $\checkmark$ $\checkmark$ 0.254 0.321 0.199 0.310
(d) $\checkmark$ $\checkmark$ $\checkmark$ 0.262 0.318 0.208 0.305
TABLE V: Quantitative ablation study was conducted to evaluate the impact of the number of hidden layers in the Adapter
Strategy Hidden layers DenseSIRST
mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
  (a) 8 0.222 0.305 0.206 0.308
(b) 16 0.243 0.318 0.214 0.310
(c) 32 0.256 0.329 0.218 0.312
(d) 64 0.266 0.336 0.230 0.323
TABLE VI: Quantitative ablation study evaluates the impact of different Feature Exchange Mechanisms within the Feature Exchange module of BAFE-Net
Strategy Exchange Method DenseSIRST
mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
  (a) Channel 0.266 0.336 0.230 0.323
(b) Spatial 0.223 0.304 0.218 0.322
(c) Channel + Spatial 0.246 0.310 0.199 0.314
(d) Spatial + Channel 0.224 0.310 0.201 0.328
TABLE VII: Ablation study on the impact of feature selection on detection performance in Channel-wise Feature Exchange
Strategy Feature Selection DenseSIRST
mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
  (a) Random Selection 0.255 0.322 0.172 0.317
(b) Fixed Selection 0.273 0.333 0.230 0.323
(c) Dynamic Selection 0.283 0.335 0.233 0.325
TABLE VIII: Ablation study on the impact of the number of exchange channels on detection performance in Top-K selection
Strategy P DenseSIRST
mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
  (a) 1/2 0.283 0.335 0.233 0.325
(b) 1/4 0.241 0.316 0.230 0.331
(c) 3/4 0.242 0.325 0.204 0.310
(d) 1/8 0.238 0.312 0.218 0.323
(e) 3/8 0.255 0.327 0.203 0.331
(f) 5/8 0.259 0.325 0.201 0.328
(g) 7/8 0.238 0.316 0.214 0.329
(h) 1/16 0.245 0.326 0.213 0.301
(i) 1/32 0.236 0.319 0.212 0.312
TABLE IX: Quantitative ablation study evaluates the impact of different backbones on BAFE-Net performance
Strategy ResNet DenseSIRST
mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
  (a) ResNet18 0.283 0.335 0.233 0.341
(b) ResNet34 0.281 0.354 0.226 0.332
(c) ResNet50 0.274 0.342 0.248 0.338
(d) ResNet101 0.261 0.317 0.241 0.336
V-C Visual Analysis
To further validate the effectiveness of our proposed methods, we conduct a visual analysis by comparing the feature activation maps of our approach with those of the top-performing competing methods in the quantitative evaluation. Fig. 6 illustrates the feature activation maps of the five most competitive methods, including our own, on representative samples from the DenseSIRST dataset.
From the visual results, we observe that our method exhibits two key advantages over the other approaches. First, the feature activation maps of our method demonstrate significantly stronger activation intensities in the target regions, indicating a higher sensitivity to the presence of infrared small targets. The deeper activation levels in the target areas suggest that our method is more effective in capturing the discriminative features of the targets, contributing to improved detection accuracy.
Second, our method generates cleaner background activation maps compared to the other methods. As evident from Fig. 6, the background regions in our activation maps exhibit consistently lower activation intensities, effectively suppressing background interference. This can be attributed to the background-aware mechanism employed by BAFE-Net, which learns to differentiate between target and background regions during the feature exchange process. By minimizing the activation levels in the background areas, our method demonstrates enhanced robustness against false positives, leading to a higher detection precision.
V-D Ablation Study
V-D1 Ablation Study on Background-Aware Gaussian Copy-Paste
We initiate our investigation from the data augmentation module, focusing on the progressive integration of our proposed Copy-Paste technique, clustering of pasted objects, Gaussian module, and Background-Aware pasting. The experimental outcomes, as outlined in Table III, generate several key insights, providing a holistic understanding of the impact of each module. Firstly, the effective enhancement of detection performance by each module is noticeable, corroborating the initial motivation of our study. The Copy-Paste technique, even when used independently, significantly boosts performance, as evidenced by the increases of 1.1% and 2.3% on the DenseSIRST dataset, when comparing strategy (a) with strategy (b).
Further refinement through clustering of pasted objects also yields substantial improvements. This is evident when comparing strategies (b) to (c), and (d) to (e), respectively, indicating the considerable value of object clustering. Applying Gaussian blur to the edges of the pasted objects, as a means to enhance their realistic appearance, also proves to be significantly beneficial in improving detection performance. This is substantiated when comparing strategies (b) with (d), and (c) with (e).
Finally, the culmination of all modules in a Background-Aware Gaussian Copy-Paste context, as depicted by strategy (f), results in further enhancements in detection capability. This is demonstrated with an additional increase of 0.8% on the DenseSIRST dataset when comparing strategy (e) with strategy (f).
V-D2 Ablation Study on the Scheme of Background-Aware Feature Exchange
The Background-Aware Feature Exchange module constitutes the core contribution of this paper, as evidenced by its significant impact on detection performance demonstrated in Table IV. To validate the effectiveness of the proposed components, we conducted an extensive ablation study by exploring various combinations of the background-based segmentation, Feature Exchange module, and adapter module.
Table IV encapsulates the results of our experimental study and yields several crucial insights. Firstly, each constituent module substantively and independently enhances the detection performance, thereby validating the underpinnings of our research. Secondly, comparison between Strategy (a) and Strategy (b) reveals a significant performance augmentation when background-based segmentation is incorporated into object detection, wherein the detection head interacts with the segmentation head. This integration accounted for an increase of 1.0% and 0.3% on the DenseSIRST dataset.
Further, juxtaposition of Strategy (b) and Strategy (c) highlights the contribution of the Feature Exchange in the interaction mode between the detection head and the segmentation head. This interaction led to a notable improvement in detection performance by 1.1% on the DenseSIRST dataset. Upon comparing Strategy (c) and Strategy (d), it becomes evident that introducing feature adapation prior to the Feature Exchange further amplifies detection performance.
As the feature adapter employs a Multi-Layer Perceptron (MLP), the number of hidden layers in the MLP is a crucial parameter that warrants careful consideration. Increasing the number of hidden layers typically improves performance but comes at the cost of increased computational complexity. To strike a balance between performance and efficiency, we conducted an experiment to determine the optimal number of hidden layers, as shown in Table  V. The experimental results demonstrate that increasing the number of hidden layers leads to a consistent improvement in detection performance. However, the performance gain becomes less significant when the number of hidden layers exceeds 64, while the parameter count increases substantially. Therefore, we opted for 64 hidden layers in the Adapter, as it provides a significant performance boost while maintaining a reasonable computational overhead compared to using 128 hidden layers or more.
V-D3 Ablation Study on Feature Exchange Mechanism
In this part, a rigorous exploration of various Feature Exchange Mechanisms was undertaken. This encompassed diverse mechanisms such as Channel Exchange and Spatial Exchange, along with a multitude of their combinations. The quest was to ascertain the most efficacious method for Feature Exchange. Channel Exchange, as a mechanism, facilitates the direct interplay at the channel level amongst different features. In contrast, Spatial Exchange is implemented by orchestrating exchanges along the spatial dimensions inherent in the feature maps. The exchange ratio was judiciously set at 0.5 to maintain a balance.
As delineated in Table VI, the superiority of the Channel Exchange method over alternative Feature Exchange mechanisms is clearly manifested. The empirical evidence underscores the robustness and optimality of Channel Exchange, thereby validating its adoption as the preferred method for Feature Exchange in our model.
V-D4 Ablation Study in Dynamic Channel Selection
In this section, we first conduct a rigorous exploration of feature selection methods. This includes random selection, fixed selection, and dynamic selection. The task is to determine the most effective feature selection method. 1) Randomly Selection: This method selects feature channels randomly, which may lead to suboptimal feature exchanges. In some cases, it could even select features that hinder the task, ultimately reducing model performance. 2) Fixed Selection: While fixed selection provides consistency during training and eliminates uncertainty, it lacks the adaptability to dynamic input features, thus limiting the model’s performance in complex or varied scenarios. 3) Dynamic Selection: dynamic selection maximizes the performance of the model and reduces unnecessary information redundancy by selecting the most useful feature channels for exchange. The exchange ratio is judiciously set to 0.5 to maintain a balance.
As shown in Table VII, the advantages of dynamic selection are obvious. Empirical evidence highlights the robustness and optimality of dynamic selection, thus confirming its adoption as the preferred method for feature selection in our model.
In addition, we conduct a rigorous exploration of the value of $p$ in Top-K selection. As shown in Table VIII, $p$ is optimal when it is 1/2.
V-D5 Impact of Backbone Depth
For infrared small target detection, the depth of the backbone network emerges as a critical factor. A shallower network, while computationally efficient, might extract features with insufficient semantic richness, thereby impinging on the overall detection performance. Conversely, an excessively deep network might overwhelm the delicate features of smaller targets amidst the background, consequently leading to a performance degradation.
We conducted a series of experiments using different ResNet architectures, varying the depth from ResNet18 to ResNet101, to identify the optimal backbone for our proposed BAFE-Net. The results of these experiments are summarized in Table IX. Notably, ResNet18 has the highest $mAP_{07}$ metric and ResNet50 has the highest $mAP_{12}$ metric.
VI Conclusion
This study tackles the key challenges in infrared small target detection by leveraging background semantics. We introduce DenseSIRST, a benchmark dataset with detailed background annotations, enabling the transition from sparse to dense target detection. Building upon this, we propose BAFE-Net, a multi-task architecture that explicitly models background semantics and facilitates cross-task feature hard-exchange. Moreover, we develop BAG-CP, a data augmentation method that selectively pastes targets into sky regions, outperforming the original Copy-Paste method. Extensive experiments validate the effectiveness of BAFE-Net and BAG-CP in improving detection accuracy and reducing false alarms. Our work underscores the importance of explicitly leveraging contextual information in infrared small target detection and paves the way for the development of more robust and context-aware algorithms in this domain.